Please help.
Determine whether the given ordered list of numbers is a solution of the system of equations:
(2, -.5)
.5x+ 8y= -3
x + 5y= -5

Answer by JulietG(1812)

.5x + 8y = -3
You're asked whether or not the pair (2,-.5) will correctly complete that equation.
Let's plug in the numbers.
.5(2) + 8(-.5) = -3
1 + -4 = -3
That one works!
.
Let's try the other.
x + 5y = -5
2 + 5(-.5) = -5
2 + -2.5 = -5
Nope. They're not equal. That one doesn't work.